OpenStudy (anonymous):

problem # 10. g(x)=log(x+5) when g(x)=5 help please

OpenStudy (amistre64):

g(x) = log(x+5) g(x) = 5 ; therefore 5 = log(x+5)

OpenStudy (cwrw238):

oopps!!

OpenStudy (cwrw238):

amis is right

OpenStudy (amistre64):

now if log means base 10 10^5 = x+5 10^5 - 5 = x

OpenStudy (anonymous):

so the answer would be x=99995. thx!

OpenStudy (anonymous):

for the same problem but g(x-3) so i started by (x-3)=log (x+5)?? is that correct how iam doing it? the professor has two answers g(x+3)=log(x+2)i dont know how he got it.

OpenStudy (amistre64):

g(x+3) is different in that it defines "x" to be a certain .... value? x = "x+3" sooo g(x+3) = log((x+3)+5) g(x+3) = log(x+8)

OpenStudy (amistre64):

it helps if I read it right to begin with

OpenStudy (amistre64):

g(x-3) is the same concept tho g(x-3) = log((x-3)+5) g(x-3) = log(x-3+5) g(x-3) = log(x+2)

OpenStudy (amistre64):

in the first bit of your post; it says g(x) = yadayada in the second bit it says x = yadayada such that: g(x) = log(x+5) g(x-3) = log((x-3)+5)

OpenStudy (anonymous):

so what are you doing is g(x-3)=log (x+5).(x-3) is that what it is? iam a little confused, sorry.